# Training Large Neural Networks With Low-Dimensional Error Feedback

### v1

@@ 1.   Introduction
In this study, we propose that high-dimensional error propagation is not strictly necessary for training deep, overparameterized networks. We introduce a novel learning rule based on Feedback Alignment that decouples forward and backward passes, enabling precise control over the dimensionality of error signals. Our method exploits the low-dimensional nature of real-world tasks to train large networks efficiently without compromising their representational capacity, achieving performance on par with backpropagation. Our theoretical analysis in linear settings highlights the potential pitfalls of naive low-dimensional error propagation and shows that a simple local online learning rule can recover backpropagation-level performance. Extending this approach to nonlinear, convolutional, and transformer-based networks, we demonstrate that learning with low-dimensional error signals is both feasible and universal. Finally, we explore how error dimensionality influences neural representations by examining its effects on receptive fields in an early visual system model, underscoring the broader implications of our findings.
@@ 2.   Background and related work
In this work, our aim is to systematically investigate how constraining the dimensionality of the error signal affects the training and performance of neural networks. To this end, we introduce a novel learning scheme, Restricted Adaptive Feedback (RAF), that allows flexible control over the dimensionality of the errors (Figure 1).
@@
    We present a novel learning rule, Restricted Adaptive Feedback (RAF), which matches BP performance while requiring minimal error signals. We provide a detailed derivation of the learning dynamics in a simple linear case, establishing a foundational understanding of how RAF operates.
@@
    We demonstrate that nonlinear networks can efficiently learn nontrivial datasets using low-dimensional error signals, highlighting the versatility of RAF in practical scenarios.
@@
    We show that more complex yet highly useful architectures, such as convolutional networks and transformers, can also be effectively trained with low-dimensional feedback.
@@
    We reveal that error dimensionality shapes the receptive fields in a model of the ventral visual system, offering new insights into the relationship between learning mechanisms and biological neural representations.
@@
Figure 1: Illustration of different approaches for propagating error to hidden layers. From left to right: Backpropagation (BP) propagates error using the exact transpose of the forward weights. Feedback Alignment (FA) replaces the transposed weights with fixed random feedback matrices, which gradually align with the forward weights during training. Restricted Adaptive Feedback (RAF) constrains error propagation through low-rank feedback matrices, limiting error dimensionality and preventing an exact mirroring of BP. Direct Restricted Adaptive Feedback (dRAF) extends RAF by enabling error signals to bypass intermediate layers, propagating directly from the output layer or other non-adjacent layers.
@@ 3.   Low-dimensional error feedback in linear networks
The goal is to learn the low-dimensional structure of $A$ from the $p$ samples using a linear neural network. For simplicity, we assume that $p$ is sufficiently large to allow the network to fully recover the structure of $A$.
@@
$\Delta W_{1}^{\mu}$ $=\eta B^{T}(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T},\text{ and }$ (4)
@@ 3.1 Learning dynamics
$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}^{T}(S-\bar{W}_{2}\bar{W}_{1}),\quad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T},$ (5)
@@
Figure 2: Learning dynamics and component alignment in linear networks. (a) Schematic of the network architecture with input dimension $n=128$, hidden layer size $k=64$, and output dimension $m=64$. The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$. (b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$, with $r=64$ and $r=8$, respectively. The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$. (d, e) Same as (b, c) but for Restricted Adaptive Feedback (RAF), where both $Q$ and $P$ are trained. (f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully. (g) Training $P$ to align with the error, as in RAF, ensures full recovery of the top $r$ singular components, improving convergence and alignment.
@@ 3.2 Training the feedback weights
Notably, the updates to the feedback weights are local and follow learning rules that were well-studied in theoretical neuroscience [18, 19, 20, 21]. Local plasticity makes our framework an attractive alternative for backpropagation in models of brain circuits.
We refer to this learning framework as Restricted Adaptive Feedback (RAF). Figure 2 compares the learning dynamics of RAF with those of BP and FA where only the $Q$ matrix is learned. The results demonstrate that RAF effectively aligns the feedback weights, enabling the network to converge to the correct solution despite the constrained error dimensionality. Furthermore, RAF will learn the top $r$ components of $\Sigma$, even if $r<d$ (Figure 2e).
@@ 4.   Nonlinear networks and complex architectures
Adapting our Restricted Adaptive Feedback (RAF) framework to nonlinear networks is straightforward because the core principles of local learning and constrained error feedback remain applicable. The local update rules in eq. (11) remain the same; the primary difference lies in the introduction of nonlinear activation functions during the propagation of signals and errors. Specifically, the forward and backward passes are modified as follows:
@@
To isolate the impact of feedback dimensionality, we applied the RAF rule in eq. (11), constraining the rank of feedback matrices in one specific layer at a time while leaving the others unrestricted. We then measured the network’s test accuracy and compared it to backpropagation (BP) as a baseline. Figure 3a shows the accuracy as a function of the feedback rank $r_{l}$, with each curve representing a different constrained layer.
@@
This compensatory effect is possible because no information is lost during the feedforward pass, unlike in a bottlenecked network. To demonstrate that the network still utilizes high-dimensional representations, we compared the performance of RAF-trained networks with constrained feedback to narrower networks without rank restrictions (Figure 3b). The results confirm that RAF-trained networks leverage their width to maintain high performance despite feedback constraints.
Moreover, as shown in Figure 3b, nonlinear networks trained with low-rank feedback in all layers using RAF can still match BP performance, demonstrating RAF’s robustness even with dimensional constraints across the entire network.
@@
Our theory shows how to propagate error signals from deeper layers to shallower ones using restricted adaptive feedback. However, error projections can also bypass intermediate layers entirely, leading to different variants of Direct Restricted Adaptive Feedback (dRAF), analogous to Direct Feedback Alignment [22]. For example, we can make direct connections from the output or the penultimate layer to earlier layers, training these connections using our algorithm (Figure 3d). As with RAF, this direct projection method matches the performance of BP when the rank of the feedback matrices satisfies $r\geq d$. This model is particularly important because it is more flexible and has greater potential to explain learning in the brain, where error signals may arrive from different pathways.
Figure 3: Restricted Adaptive Feedback (RAF) efficiently trains nonlinear networks on CIFAR-10. (a) Constraining feedback in any layer to $r=d=10$ matches full BP performance. Different curves ($B_{1}$, $B_{2}$, and $B_{3}$) correspond to restricting feedback at different layers. (b) Reducing feedback dimensionality has minimal impact on performance, whereas reducing network width significantly degrades accuracy, indicating that high-dimensional representations are still utilized. (c) Minimal error dimensionality is sufficient for learning, as shown by subsampling classes from CIFAR-100. All feedback matrices are constrained to the same rank $r$. (d) Direct RAF: all layers receive a low-dimensional error either directly from the output layer (solid) or from the penultimate layer (dashed), both converging to BP-level performance. (e) Performance of a 4-block VGG-like convolutional network trained with RAF, constraining the layers with 512 channels. Inset: Training the same network with all layers constrained to a fraction of their size; the smallest layer has 64 channels. (f) A transformer-based architecture trained to classify images, demonstrating that RAF generalizes to more advanced network architectures. See Supplemental Information for details on all implementations.
@@
Our previous results demonstrate that fully connected networks can learn effectively from minimal error signals, even on complex datasets, matching backpropagation (BP) performance. Here, we extend this investigation to convolutional architectures.
Training convolutional networks with Feedback Alignment (FA) is notoriously difficult [8, 23]. Recent work has made progress by learning feedback weights within the FA framework [24], but our approach differs by restricting the error signal’s dimensionality using Restricted Adaptive Feedback (RAF). We aim to determine whether convolutional networks can also benefit from low-dimensional error feedback.
We trained a VGG-like convolutional network [25] with four blocks and batch normalization on the CIFAR-10 dataset (see Appendix for details). Using RAF, we decoupled the error propagation from the feedforward pass in all layers. Initially, we constrained the feedback error only in the blocks containing 512 (Figure 3e). Consistent with our findings in fully connected networks, convolutional networks learn well with a feedback matrix with a rank similar to the number of classes, $d=10$.
@@
To investigate this question, we applied our approach to visual transformer architectures trained for image recognition on the CIFAR-10 dataset. Each weight matrix in the transformer architecture, including key, query, and value matrices, was trained using low-dimensional error feedback from the subsequent layer. Feedback matrices were trained using the RAF framework (see the appendix for details). The network’s performance as a function of the feedback rank, which was kept identical for all matrices, is shown in Figure 3f.
Our results indicate that, consistent with our findings in both nonlinear and convolutional networks, an error signal with dimensionality as low as the task dimensionality (i.e. the number of output classes) is sufficient to train transformer architectures effectively (Figure 3f). Despite the intricate nature of self-attention computations, training with low-dimensional feedback yielded performance comparable to conventional backpropagation on CIFAR-10 with rank $r=10$. Notably, training transformers with low-dimensional feedback using RAF not only matched but, in some cases, consistently outperformed backpropagation by a small margin. This improvement was observed even when hyperparameters were independently optimized for both BP and RAF. The improved generalization could be understood as a regularization effect achieved by constraining the feedback. However, since we observe it only in the transformer architecture, it is not a general property of a low-dimensional error signal, and we leave further investigation to a future study.
Figure 4: Receptive fields (RFs) are shaped by feedback dimensionality rather than feedforward constraints. (a) Model of the early ventral visual stream. Top: Retinal-to-cortical connectivity is modeled as a bottleneck in the network; adapted from [29]. Middle: A bottleneck is applied only to the feedback connections, restricting error dimensionality. Bottom: Using direct RAF (dRAF), retinal convolutional layers receive error signals from the penultimate layer, enabling full error feedback even with a feedforward bottleneck. (b) Receptive fields (RFs) in a model trained with BP, corresponding to the setup in (a) top. Left: No bottleneck. Right: Bottleneck at the retinal output, which induces center-surround RFs. (c) Training the retinal output with RAF. Restricting the rank of the feedback matrix produces center-surround RFs. (d) RFs in a network with a bottleneck at the retinal output but full-dimensional error feedback via dRAF. The presence of high-dimensional error signals results in oriented RFs.
@@ 5.   Error dimensionality shapes neural receptive fields
Lindsey et al. found that in convolutional models of the visual system, narrow feedforward bottlenecks between the retina and the brain led to center-surround receptive fields in the retinal layer, similar to mammals [29]. Wider bottlenecks resulted in orientation-selective receptive fields, as seen in salamanders. We hypothesize that these effects are due to constraints on the error signal reaching the retina, rather than the feedforward bottlenecks themselves.
To test this hypothesis, we trained a model similar to that of Lindsey et. al  [29] but with full-width layers throughout. Instead of constraining the feedforward pathway, we constrained only the error signal using a low-rank feedback matrix trained with RAF (Figure 4c), thereby isolating the impact of error dimensionality on neural representations. We extracted the receptive fields of neurons in the retinal layer using visualization techniques [30] (Figure 4d).
Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer. To further validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dRAF to train the retinal layer without restricting the backward pathway. In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix).
@@ Appendix A Full linear theory
in this section, we provide the analysis of the linear networks for both Feedback Alignment and RAF
@@
Full derivation of Restricted Adaptive Feedback (RAF) learning dynamics

### v2

@@ Abstract
Significance statement
Conventional wisdom holds that training large neural networks requires propagating precise, high-dimensional error gradients through every neuron, a process with limited biological plausibility. Our work demonstrates that neural networks can be trained effectively using low-dimensional error signals, comparable to the inherent complexity of the task rather than the size of the network. By developing a novel, biologically plausible learning paradigm based on adaptive low-dimensional feedback, we show theoretically and empirically that minimal error dimensionality can achieve competitive performance even in complex neural architectures. These findings challenge existing assumptions about neural optimization and suggest fundamental constraints that may shape learning in both artificial and biological systems.
@@ 1.   Introduction
In this study, we propose that high-dimensional error propagation is not strictly necessary for training deep, overparameterized networks. We introduce a novel learning rule based on Feedback Alignment that decouples forward and backward passes, enabling precise control over the dimensionality of error signals. Our method exploits the low-dimensional nature of real-world tasks to train large networks efficiently without compromising their representational capacity, achieving performance on par with backpropagation. A theoretical analysis in linear settings highlights the potential pitfalls of naive low-dimensional error propagation and shows that a simple local online learning rule can recover backpropagation-level performance. Extending this approach to nonlinear, convolutional, and transformer-based networks, we demonstrate that learning with low-dimensional error signals is feasible and universal. Finally, we explore how error dimensionality influences neural representations by examining its effects on receptive fields in an early visual system model, underscoring the broader implications of our findings.
@@ 2.   Background and related work
In this work, our objective is to systematically investigate how constraining the dimensionality of the error signal affects the training and performance of large neural networks. To this end, we introduce a novel learning scheme that enables deep layers to be learned using low-rank feedback matrices, which allow flexible control over the dimensionality of the loss gradients in each layer (Figure 1).
@@
    We present a novel local learning rule for low-rank feedback weights, which matches BP performance while requiring minimal error signals. We provide a detailed derivation of the learning dynamics in a simple linear case, establishing the foundation for why and how the feedback weights should be learned.
@@
    We demonstrate that large and deep nonlinear networks can efficiently learn nontrivial datasets using low-dimensional error signals and highlight that the feedback dimensionality should scale with the task rather than the size of the network.
@@
    We successfully train complex yet highly useful architectures, such as convolutional networks and visual transformers, using low-dimensional feedback.
@@
    We study a simplified model of the ventral visual system and show that the emerging receptive fields depend on the error-feedback dimensionality. These findings offer new insights into the relationship between learning mechanisms and biological neural representations.
@@
Figure 1: Illustration of different approaches for propagating error to hidden layers. From left to right: Backpropagation (BP) propagates error using the exact transpose of the forward weights. Feedback Alignment (FA) replaces the transposed weights with fixed random feedback matrices, which gradually align with the forward weights during training. Our work framework provides low-dimensional feedback (LDF) using low-rank matrices, allowing full control of the dimnsionality of the gradients in each layer. Finally, direct low-dimensional feedback (dLDF) extends our LDF architecture by enabling error signals to bypass intermediate layers, propagating directly from the output layer or other non-adjacent layers.
@@ 3.   Low-dimensional error feedback in linear networks
The goal is to learn the low-dimensional structure of $A$ from the $p$ samples using a linear neural network. For simplicity, we assume that $p$ is large enough to allow the network to fully recover the structure of $A$.
@@
$\Delta W_{1}^{\mu}$ $=\eta B(\boldsymbol{y}^{\mu}-W_{2}W_{1}\boldsymbol{x}^{\mu})\boldsymbol{x}^{\mu T},\text{ and }$ (4)
@@ 3.1 Learning dynamics
$\tau\frac{d\bar{W}_{1}}{dt}=\bar{B}(S-\bar{W}_{2}\bar{W}_{1}),\quad\tau\frac{d\bar{W}_{2}}{dt}=(S-\bar{W}_{2}\bar{W}_{1})\bar{W}_{1}^{T},$ (5)
@@
Figure 2: Learning dynamics and component alignment in linear networks. (a) Schematic of the network architecture with input dimension $n=128$, hidden layer size $k=64$, and output dimension $m=64$. The feedback matrix $B$ is factorized as $B=QP$ and constrained to rank $r$. (b, c) Comparison of theoretical predictions (dashed) and numerical simulations (solid) for low-rank Feedback Alignment (FA), training only $Q$[10], with $r=64$ and $r=8$, respectively. The $y$-axis shows the alignment of singular vectors between $W_{2}W_{1}$ and $\Sigma_{io}$. (d, e) Same as (b, c) but training both and $Q$ and $P$ using Eqs (7) and (8) . (f) When $P$ is not trained, the singular modes align on average (bold), but the top $r$ components fail to recover fully. (g) Training $P$ to align with the error ensures full recovery of the top $r$ singular components, improving convergence and alignment.
@@ 3.2 Training the feedback weights
Notably, the updates to the feedback weights are local and follow learning rules that were well-studied in theoretical neuroscience [18, 19, 20, 21]. Local plasticity makes our framework an attractive alternative for backpropagation in brain circuit models.
We refer to this learning framework as low-dimensional feedback alignment (LDFA). Figure 2 compares the learning dynamics of LDFA with those of BP and FA where only the $Q$ matrix is learned. The results demonstrate that LDFA effectively aligns the feedback weights, enabling the network to converge to the correct solution despite the constrained error dimensionality. Furthermore, LDFA will learn the top $r$ components of $\Sigma$, even if $r<d$ (Figure 2e).
@@ 4.   Nonlinear networks and complex architectures
Adapting our low-dimensional feedback alignment (LDFA) framework to nonlinear networks is straightforward because the core principles of local learning and constrained error feedback remain applicable. The local update rules in eq. (11) remain the same; the primary difference lies in the introduction of nonlinear activation functions during the propagation of signals and errors. Specifically, the forward and backward passes are modified as follows:
@@
To isolate the impact of feedback dimensionality, we applied the LDFA rule in eq. (11), constraining the rank of feedback matrices in one specific layer at a time while leaving the others unrestricted. We then measured the network’s test accuracy and compared it to backpropagation (BP) as a baseline. Figure 3a shows the accuracy as a function of the feedback rank $r_{l}$, with each curve representing a different constrained layer.
@@
This compensatory effect is possible because no information is lost during the feedforward pass, unlike in a bottlenecked network. To demonstrate that the network still utilizes high-dimensional representations, we compared the performance of LDFA-trained networks with constrained feedback to narrower networks without rank restrictions (Figure 3b). The results confirm that LDFA-trained networks leverage their width to maintain high performance despite feedback constraints.
Moreover, as shown in Figure 3b, nonlinear networks trained with low-rank feedback in all layers using LDFA can still match BP performance, demonstrating LDFA’s robustness even with dimensional constraints across the entire network.
@@
Our theory shows how to propagate error signals from deeper layers to shallower ones using restricted adaptive feedback. However, error projections can also bypass intermediate layers entirely, leading to different variants of direct low-dimensional feedbackalignment (dLDFA), analogous to Direct Feedback Alignment [22]. For example, we can make direct connections from the output or the penultimate layer to earlier layers, training these connections using our algorithm (Figure 3d). As with LDFA, this direct projection method matches the performance of BP when the rank of the feedback matrices satisfies $r\geq d$. This model is particularly important because it is more flexible and has greater potential to explain learning in the brain, where error signals may arrive from different pathways.
Figure 3: Low-dimensional feedback alignment (LDFA) efficiently trains nonlinear networks on CIFAR-10. (a) Constraining feedback in any layer to $r=d=10$ matches full BP performance. Different curves ($B_{1}$, $B_{2}$, and $B_{3}$) correspond to restricting feedback at different layers. (b) Reducing feedback dimensionality has minimal impact on performance, whereas reducing network width significantly degrades accuracy, indicating that high-dimensional representations are still utilized. (c) Minimal error dimensionality is sufficient for learning, as shown by subsampling classes from CIFAR-100. All feedback matrices are constrained to the same rank $r$. (d) Direct LDFA: all layers receive a low-dimensional error either directly from the output layer (solid) or from the penultimate layer (dashed), both converging to BP-level performance. (e) Performance of a 4-block VGG-like convolutional network trained with LDFA, constraining the layers with 512 channels. Inset: Training the same network with all layers constrained to a fraction of their size; the smallest layer has 64 channels. (f) A transformer-based architecture trained to classify images, demonstrating that LDFA generalizes to more advanced network architectures. See Supplemental Information for details on all implementations.
@@
Our previous results demonstrate that fully connected networks can learn effectively from minimal error signals, even with complex datasets, matching backpropagation (BP) performance. Here, we extend this investigation to convolutional architectures.
Training convolutional networks with Feedback Alignment (FA) is notoriously difficult [8, 23]. Recent work has made progress by learning feedback weights within the FA framework [24], but our approach differs by restricting the error signal’s dimensionality using LDFA. We aim to determine whether convolutional networks can also benefit from low-dimensional error feedback.
We trained a VGG-like convolutional network [25] with four blocks and batch normalization on the CIFAR-10 dataset (see the appendix for details). Using LDFA, we decoupled error propagation from the feedforward pass in all layers. Initially, we constrained the feedback error only in the blocks containing 512 (Figure 3e). Consistent with our findings in fully connected networks, convolutional networks learn well with a feedback matrix with a rank similar to the number of classes, $d=10$.
@@
To investigate this question, we applied our approach to visual transformer architectures trained for image recognition on the CIFAR-10 dataset. Each weight matrix in the transformer architecture, including key, query, and value matrices, was trained using low-dimensional error feedback from the subsequent layer. Feedback matrices were trained using the LDFA framework (see the appendix for details). The network’s performance as a function of the feedback rank, which was kept identical for all matrices, is shown in Figure 3f.
Our results indicate that, consistent with our findings in both nonlinear and convolutional networks, an error signal with dimensionality as low as the task dimensionality (i.e. the number of output classes) is sufficient to train transformer architectures effectively (Figure 3f). Despite the intricate nature of self-attention computations, training with low-dimensional feedback yielded performance comparable to conventional backpropagation on CIFAR-10 with rank $r=10$. Notably, training transformers with low-dimensional feedback using LDFA not only matched but, in some cases, consistently outperformed backpropagation by a small margin. This improvement was observed even when hyperparameters were independently optimized for both BP and LDFA. The improved generalization could be understood as a regularization effect achieved by constraining the feedback. However, since we observe it only in the transformer architecture, it is not a general property of a low-dimensional error signal, and we leave further investigation to a future study.
Figure 4: Receptive fields (RFs) are shaped by feedback dimensionality rather than feedforward constraints. (a) Model of the early ventral visual stream. Top: Retinal-to-cortical connectivity is modeled as a bottleneck in the network; adapted from [29]. Middle: A bottleneck is applied only to the feedback connections, restricting error dimensionality. Bottom: Using direct LDFA (dLDFA), retinal convolutional layers receive error signals from the penultimate layer, enabling full error feedback even with a feedforward bottleneck. (b) Receptive fields (RFs) in a model trained with BP, corresponding to the setup in (a) top. Left: No bottleneck. Right: Bottleneck at the retinal output, which induces center-surround RFs. (c) Training the retinal output with LDFA. Restricting the rank of the feedback matrix produces center-surround RFs. (d) RFs in a network with a bottleneck at the retinal output but full-dimensional error feedback via dLDFA. The presence of high-dimensional error signals results in oriented RFs.
@@ 5.   Error dimensionality shapes neural receptive fields
Lindsey et al. found that in convolutional models of the visual system, narrow feedforward bottlenecks between the retina and the brain led to center-surround receptive fields in the retinal layer, similar to mammals [29]. Wider bottlenecks resulted in orientation-selective receptive fields, as seen in salamanders. We hypothesize that these effects are due to constraints on the error signal reaching the retina rather than the feedforward bottlenecks themselves.
To test this hypothesis, we trained a model similar to that of Lindsey et al.  [29], but with full-width layers throughout. Instead of constraining the feedforward pathway, we constrained only the error signal using a low-rank feedback matrix trained with LDFA (Figure 4c), thereby isolating the impact of error dimensionality on neural representations. We extracted the receptive fields of neurons in the retinal layer using visualization techniques [30] (Figure 4d).
Consistent with our hypothesis, constraining the feedback rank led to the emergence of center-surround receptive fields in the retinal layer. To validate our hypothesis, we trained a model that included the feedforward bottleneck, similar to [29] but used dLDFA to train the retinal layer without restricting the backward pathway. In line with our expectations, the retinal receptive fields exhibited orientation selectivity (see the appendix).
@@ Appendix A Full linear theory
in this section, we provide the analysis of the linear networks for both Feedback Alignment and LDFA
@@
Full derivation of low-dimensional feedback alignment (LDFA) learning dynamics